OpenStudy (anonymous):

if you weight 480N on eath how much do you weight on the moon

OpenStudy (anonymous):

gravity is only 1/6 the strength on the moon as on Earth. So if 'g' for Earth is 9.8 m/s^2, then on the moon it is 9.8/6 = 1.63 m/s^2 If you weight 480 N on Earth: 480 = mg = m * 9.8 --> m = 48.98 kg So on the moon: W = 48.98 * 1.63 = 79.84 N

OpenStudy (michele_laino):

please note that gravity on the Moon is g/6, where g=9.81 m/sec^2